Arlington LLC purchased an automobile for $80,000 on July 5, 2018. What is Arlington's depreciation expense for 2018 if its business use percentage is 80 percent (ignore any possible bonus depreciation)? (Use MACRS Table 1, and Exhibit 10-10.)

Multiple Choice

$6,125.

$8,100.

$8,000.

$12,000.

None of the choices are correct.

Bateman Corporation sold an office building that it used in its business for $800,850. Bateman bought the building ten years ago for $599,575 and has claimed $201,275 of depreciation expense. What is the amount and character of Bateman's gain or loss?

Multiple Choice

$40,255 ordinary and $362,295 §1231 gain.

$201,275 ordinary and $201,275 §1231 gain.

$402,550 ordinary gain.

$402,550 capital gain.

None of the choices are correct.
Answer #1

1) Considering calendar year as Jan-Dec and depreciation rate as 25% (as the same is not mentioned)

As the asset was purchased in july i.e after 6 months of start of FY, thus depreciation rate will behalf i.e 12.5%. Further, as the asset was used only for 80% of time for business. Depreciation amount will be as below:-

80,000*12.5%*80%= $8000

2) Sale price of office building- $800,850

Purchase price 10 years ago was-$599,575,

Depreciation expense- $201,275

Present cost price of building- Purchase price- Depreciation till date= 599,575- 201,275= $398,300

As the asset is a capital asset, the sale of the same will generate capital gain of- sales price-present cost=800,850-398,300=$402,550